OpenStudy (hijacktrolo):

Choose the point-slope form of the equation below that represents the line that passes through the point (6, −3) and has a slope of one half. y − 6 = one half(x + 3) y = one halfx − 6 y + 3 = one half(x − 6) x − 2y = 12

OpenStudy (hijacktrolo):

please help

OpenStudy (will.h):

We will have to use the following formula y - y1 = m(x - x1) M is the slope and we know we have the point (6,-3) so the y1 is -3 and x1 is 6 and now substitute y - (-3) = 1/2 (x - 6) Simplify the left side y + 3 = 1/2 (x-6)

OpenStudy (hijacktrolo):

thanks